# Soft-Label Caching and Sharpening for Communication-Efficient Federated Distillation

### v1

@@ IV-A Experimental Setup
Datasets and Partitioning: SCARLET was evaluated using three combinations of private and public datasets for image classification tasks under heterogeneous data conditions. The detailed dataset configurations are summarized in Table II.
@@
To ensure the robustness of SCARLET in practical scenarios, we deliberately avoided methods that split a single dataset into public and private datasets, such as in [3]. Instead, we adopted the approach inspired by [22], where separate datasets with no overlap in image content are used for public and private datasets. This choice reflects real-world conditions, as acquiring a public dataset with identical distributions to private data is often impractical. By utilizing distinct datasets, SCARLET aligns more closely with application contexts where such separations are necessary, enhancing the realism and relevance of the evaluation.
The private and public datasets used in our experiments are entirely distinct, with no direct class overlaps. Specifically, CIFAR-10 and CIFAR-100 [27], despite being from the same dataset family, have no overlapping classes. For instance, CIFAR-10 includes categories such as “automobile” and “truck,” while CIFAR-100 contains related but separate categories like “bus” and “pickup truck.” Similarly, Tiny ImageNet [28] and Caltech-256 [29] cover broad object categories without any direct overlap in image content. While some semantic similarities exist between certain categories across datasets, they do not affect the integrity of the experimental setup.
Data heterogeneity across clients was simulated using a Dirichlet distribution [30] with varying $\alpha$ values. Larger $\alpha$ values approximate IID distributions, while smaller values result in each client holding data from fewer classes. Fig. 5 visualizes the non-IID sample distributions among clients. Smaller $\alpha$ values lead to skewed distributions, where each client primarily holds data from fewer classes. This setup evaluates SCARLET’s performance under varying degrees of data heterogeneity and better simulates challenging FL environments.
@@
Figure 5: Visualization of private dataset distributions across clients under different Dirichlet parameters ($\alpha$). Smaller $\alpha$ values result in more heterogeneous distributions, where clients predominantly hold data from fewer classes.
Models: Table III outlines the model architectures and input image sizes used for each private dataset. For image classification tasks, we tailored ResNet [31] architectures to match each dataset’s characteristics. CIFAR-10 and CIFAR-100 both use $32\times 32$ images; however, the increased complexity of CIFAR-100, with its 100 classes, necessitates a deeper ResNet-32 model. For Tiny ImageNet, which has $64\times 64$ images, we used a ResNet-18, which is well suited for intermediate-resolution inputs.
@@
Comparison Methods and Hyperparameter Settings: We evaluated SCARLET’s communication efficiency and model accuracy against four state-of-the-art methods: DS-FL [3], CFD [4], COMET [22], and Selective-FD [23].
Unless otherwise mentioned, the following hyperparameter settings were common across all methods. The number of clients was set to 100, each performing 5 local epochs per training round. The learning rate for both local client training and client-server distillation was set to 0.1, following [3]. Each training round utilized 1,000 unlabeled public data samples. For simplicity, we employed stochastic gradient descent (SGD) without momentum or weight decay as the optimizer. The number of communication rounds was determined to ensure sufficient convergence for each private dataset setting: we used 2000 rounds for CIFAR-10 ($\alpha=0.05$), 3000 rounds for CIFAR-10 ($\alpha=0.3$) and CIFAR-100, and 1000 rounds for Tiny ImageNet.
In contrast, method-specific hyperparameters critical to performance, such as the ERA temperature $T$ in DS-FL, were individually optimized through hyperparameter tuning. Detailed hyperparameter configurations for each comparative method are provided in Appendix B.
Evaluation Metrics: The evaluation of SCARLET is based on two primary metrics: cumulative communication cost and test accuracy. The cumulative communication cost is defined as the total amount of data exchanged between the server and clients, including soft-labels and other parameters, summed over all clients across communication rounds. We exclude the one-time cost for retrieving the public dataset from this calculation, as it is common to all compared methods and does not affect relative performance.
@@
Figure 6: Comparison of each client’s private dataset and test dataset distributions. Both are sampled using the same Dirichlet distribution, ensuring that each client’s test set closely matches the class proportions of its private dataset. This setup enables personalized, non-IID evaluation at the client side.
@@ IV-B Experimental Results
Centralized Learning (Non-FL) Reference Results: Table IV shows the performance obtained by centralized learning, where all clients upload their labeled private data to a central server, and a single global model is trained. These results serve as a reference to illustrate achievable performance levels without the constraints of FL. In this setting, since all data is aggregated at a central location, the resulting training data is effectively IID. This often leads to significantly higher accuracy compared to FL conducted under non-IID data conditions.
@@
Tiny-ImageNet ResNet-18 39.5%
@@
Figure 7: Test accuracy versus communication cost for SCARLET, DS-FL (baseline), and SOTA methods (CFD, COMET, Selective-FD) on CIFAR-10 under two non-IID conditions: $\alpha=0.05$ (strong non-IID) and $\alpha=0.3$ (weaker non-IID).
@@
Comparative Evaluation: We evaluate SCARLET’s communication efficiency and model accuracy by comparing it against four state-of-the-art methods—DS-FL [3], CFD [4], COMET [22], and Selective-FD [23]—as well as an Individual baseline that represents isolated client training without collaboration. All results were averaged across three random seeds, with details of hyperparameter settings for comparative methods provided in Appendix B.
@@
In all experiments, SCARLET employed a soft-label caching mechanism with a cache duration of $D=50$ and utilized the Enhanced ERA strategy to adapt to non-IID data distributions. The sharpness parameter $\beta$ in Enhanced ERA was tuned to reflect the degree of data heterogeneity. For CIFAR-10, $\beta$ was set to $2.0$ under strongly non-IID conditions ($\alpha=0.05$) and $1.0$ under less heterogeneous conditions ($\alpha=0.3$). For CIFAR-100 and Tiny ImageNet, $\beta=1.25$ yielded optimal results. A more detailed evaluation of the effects of varying $\alpha$ and $\beta$ is provided in the next section.
@@
Fig. 7 presents the relationship between server-side and client-side test accuracy and cumulative communication cost on CIFAR-10 under strong and moderate non-IID settings. SCARLET outperformed other methods under strongly non-IID conditions ($\alpha=0.05$), achieving approximately 2% higher server-side accuracy while reducing communication costs by over 50%. On the client side, SCARLET matched the peak accuracy of Selective-FD but with similar communication savings. This highlights the effectiveness of SCARLET’s caching mechanism and adaptive aggregation in mitigating challenges posed by heterogeneous data distributions.
@@
Under less severe non-IID conditions ($\alpha=0.3$), SCARLET maintained server-side accuracy comparable to the best-performing method, Selective-FD, while achieving over 50% lower communication cost. However, Selective-FD demonstrated superior client-side convergence speed and final accuracy, underscoring its strength in personalized learning scenarios. In contrast, DS-FL struggled to achieve high accuracy under these conditions, regardless of how the temperature $T$ in its ERA mechanism was tuned. This limitation can be traced to the presence of aggregated soft-labels with inherently low entropy. When $T$ is reduced excessively, these low-entropy soft-labels approach a one-hot representation, diminishing the effectiveness of the distillation process. Conversely, increasing $T$ too much causes a sudden spike in entropy for certain soft-labels, exacerbating inconsistencies in entropy reduction across public datasets. These issues result in both lower accuracy and performance degradation over time, likely due to overfitting.
@@
(a) Server (CIFAR-100)
(b) Client (CIFAR-100)
(c) Server (Tiny ImageNet)
(d) Client (Tiny ImageNet)
Figure 8: Test accuracy versus communication cost results on CIFAR-100 and Tiny ImageNet for SCARLET and SOTA methods (DSFL, CFD, COMET, Selective-FD). These results are still under investigation.
Fig. 8 presents the results for CIFAR-100 and Tiny ImageNet, demonstrating the scalability and robustness of SCARLET. The soft-label caching mechanism and Enhanced ERA consistently achieved high communication efficiency and competitive accuracy across datasets with more complex label distributions and greater heterogeneity.
@@
Effectiveness of the Soft-Label Caching Mechanism in Other SOTA Methods: The soft-label caching mechanism, a core component of SCARLET, is designed to alleviate communication inefficiencies in distillation-based FL. Since such inefficiencies are not specific to SCARLET, the mechanism can also be applied to other SOTA methods to improve their communication efficiency. To evaluate its broader impact, we integrated the caching mechanism into three representative FL methods: CFD, COMET, and Selective-FD. DS-FL was excluded from this evaluation, as it becomes conceptually similar to SCARLET when caching is introduced. For all methods, we used a conservative cache duration of $D=25$ to balance communication reduction with the risk of overfitting to stale predictions.
(a) Server (CFD)
(b) Client (CFD)
(c) Server (COMET)
(d) Client (COMET)
(e) Server (Selective-FD)
(f) Client (Selective-FD)
Figure 9: Impact of the soft-label cache mechanism integration on test accuracy and communication cost for SCARLET applied to CFD, COMET, and Selective-FD. Plots for both server and client sides demonstrate significant communication cost reductions without substantial accuracy degradation.
@@
Fig. 9 presents the server-side and client-side test accuracy versus cumulative communication cost for each method with the soft-label caching mechanism applied. The results show that incorporating the caching mechanism reduced server-side communication costs by approximately 50% across all three methods, without substantial loss in target model accuracy. Notably, CFD and Selective-FD maintained accuracy levels comparable to their original implementations, indicating that the mechanism imposes minimal performance overhead. On the client side, although communication reductions were less dramatic due to the quicker convergence of personalized models, notable gains were still observed. Selective-FD, in particular, preserved strong personalized accuracy while benefiting from improved communication efficiency. COMET showed slight accuracy degradation in later training rounds, a trend also reflected in its server-side performance.
@@
These findings demonstrate the generalizability and effectiveness of the soft-label caching mechanism. Its seamless integration into a variety of distillation-based FL methods—including CFD, COMET, and Selective-FD—highlights its potential as a practical, drop-in component for enhancing communication efficiency. By incorporating this mechanism, existing FL frameworks can significantly reduce communication overhead while maintaining strong global and personalized model performance, thereby contributing to the scalability and practicality of FL systems.

### v3

@@ IV-A Experimental Setup
IV-A1 Datasets and Partitioning
SCARLET was evaluated using three combinations of private and public datasets for image classification tasks under heterogeneous data conditions. The detailed dataset configurations are summarized in Table II.
@@ IV-A1 Datasets and Partitioning
To ensure the robustness of SCARLET in practical scenarios, we deliberately avoided methods that split a single dataset into public and private datasets, such as in [dsfl]. Instead, we adopted the approach inspired by [comet], where separate datasets with no overlap in image content are used for public and private datasets. This choice reflects real-world conditions, as acquiring a public dataset with identical distributions to private data is often impractical. By utilizing distinct datasets, SCARLET aligns more closely with application contexts where such separations are necessary, enhancing the realism and relevance of the evaluation.
The private and public datasets used in our experiments are entirely distinct, with no direct class overlaps. Specifically, CIFAR-10 and CIFAR-100 [cifar], despite being from the same dataset family, have no overlapping classes. For instance, CIFAR-10 includes categories such as “automobile” and “truck” while CIFAR-100 contains related but separate categories like “bus” and “pickup truck.” Similarly, Tiny ImageNet [tinyimagenet] and Caltech-256 [caltech] cover broad object categories without any direct overlap in image content. While some semantic similarities exist between certain categories across datasets, they do not affect the integrity of the experimental setup.
Data heterogeneity across clients was simulated using a Dirichlet distribution [dirichlet] with varying $\alpha$ values. Larger $\alpha$ values approximate IID distributions, while smaller values result in each client holding data from fewer classes. Fig. 6 visualizes the non-IID sample distributions among clients. Smaller $\alpha$ values lead to skewed distributions, where each client primarily holds data from fewer classes. This setup evaluates SCARLET’s performance under varying degrees of data heterogeneity and better simulates challenging FL environments.
@@
Figure 6: Visualization of private dataset distributions across clients under different Dirichlet parameters ($\alpha$). Smaller $\alpha$ values result in more heterogeneous distributions, where clients predominantly hold data from fewer classes. (6a) and (6b) show CIFAR-10 with varying $\alpha$, while (6c) and (6d) illustrate CIFAR-100 and Tiny ImageNet, respectively.
IV-A2 Models
Table III outlines the model architectures and input image sizes used for each private dataset. For image classification tasks, we tailored ResNet [resnet] architectures to match each dataset’s characteristics. CIFAR-10 and CIFAR-100 both use $32\times 32$ images; however, the increased complexity of CIFAR-100, with its 100 classes, necessitates a deeper ResNet-32 model. For Tiny ImageNet, which has $64\times 64$ images, we used a ResNet-18, which is well suited for intermediate-resolution inputs.
@@ IV-A2 Models
IV-A3 Comparison Methods and Hyperparameter Settings
We evaluated SCARLET’s communication efficiency and model accuracy against four state-of-the-art methods: DS-FL [dsfl], CFD [cfd], COMET [comet], and Selective-FD [selectivefd].
Unless otherwise mentioned, the following hyperparameter settings were common across all methods. The number of clients was set to 100, each performing 5 local epochs per training round. The learning rate for both local client training and client-server distillation was set to 0.1, following [dsfl]. Each training round utilized 1,000 unlabeled public data samples. For simplicity, we employed stochastic gradient descent (SGD) without momentum or weight decay as the optimizer.
The total number of communication rounds was determined to ensure our proposed method, SCARLET, and most of the baselines reached a stable performance plateau where its accuracy saturated. While this experimental setup uses a global test set for analysis, we validate in Appendix D that this convergence plateau is observable in practice using accessible proxy metrics (e.g., local validation loss) that do not require such a test set. Accordingly, unless otherwise mentioned, we used 3000 rounds for CIFAR-10 and CIFAR-100, and 1000 rounds for Tiny ImageNet. This duration was deemed sufficient to provide a fair comparison of the convergence speed and accuracy against baseline methods within this operational window.
In contrast, method-specific hyperparameters critical to performance, such as the ERA temperature $T$ in DS-FL, were individually optimized through hyperparameter tuning. Detailed hyperparameter configurations for each comparative method are provided in Appendix E.
@@ IV-A3 Comparison Methods and Hyperparameter Settings
Figure 7: Comparison of each client’s private and test dataset distributions. Both datasets are generated using the same Dirichlet distribution, ensuring that each client’s test set maintains class proportions similar to its private dataset. This configuration supports personalized and non-IID evaluation on the client side.
IV-A4 Evaluation Metrics
The evaluation of SCARLET is based on two primary metrics: cumulative communication cost and test accuracy. The cumulative communication cost is defined as the total amount of data exchanged between the server and clients, including soft-labels and other parameters, summed over all clients across communication rounds. We exclude the one-time cost for retrieving the public dataset from this calculation, as it is common to all compared methods and does not affect relative performance. Furthermore, to provide a deeper analysis that reflects real-world network conditions with often asymmetric bandwidth, we also present a detailed breakdown of the uplink (client-to-server) and downlink (server-to-client) communication costs for our main comparison.
@@ IV-B Experimental Results
Tiny ImageNet ResNet-18 39.5%
IV-B1 Centralized Learning (Non-FL) Reference Results
Table IV shows the performance obtained by centralized learning, where all clients upload their labeled private data to a central server, and a single global model is trained. These results serve as a reference to illustrate achievable performance levels without the constraints of FL. In this setting, since all data is aggregated at a central location, the resulting training data is effectively IID. This often leads to significantly higher accuracy compared to FL conducted under non-IID data conditions.
@@ IV-B1 Centralized Learning (Non-FL) Reference Results
Figure 8: Test accuracy vs. communication cost for SCARLET and other state-of-the-art methods on CIFAR-10. SCARLET shows the largest advantage under strong non-IID conditions (8a, 8b). This improvement results from the synergy of our two main contributions: the soft-label caching mechanism reduces data transmission, while the Enhanced ERA mechanism effectively mitigates the negative impact of data heterogeneity. Under moderate non-IID conditions (8c, 8d), SCARLET maintains competitive accuracy with similar communication efficiency.
IV-B2 Comparative Evaluation
@@
We evaluate SCARLET’s communication efficiency and model accuracy by comparing it against four state-of-the-art methods—DS-FL [dsfl], CFD [cfd], COMET [comet], and Selective-FD [selectivefd]—as well as an Individual baseline that represents isolated client training without collaboration. All results were averaged across three random seeds, with details of hyperparameter settings for comparative methods provided in Appendix E.
@@
In all experiments, SCARLET employed a soft-label caching mechanism with a cache duration of $D=50$ and utilized the Enhanced ERA strategy to adapt to non-IID data distributions. The sharpness parameter $\beta$ in Enhanced ERA was tuned to reflect the degree of data heterogeneity. For CIFAR-10, $\beta$ was set to $1.5$ under strongly non-IID conditions ($\alpha=0.05$) and $1.0$ under moderate non-IID conditions ($\alpha=0.3$). For CIFAR-100 and Tiny ImageNet, $\beta=1.25$ was used. These parameters were chosen as they represent a strong balance of performance and efficiency, as demonstrated in our detailed ablation studies on cache duration $D$ (Section IV-B4) and aggregation sharpness $\beta$ (Section IV-B5).
@@
Fig. 8 presents the relationship between server-side and client-side test accuracy and cumulative communication cost on CIFAR-10 under strong and moderate non-IID settings. SCARLET outperformed other methods under strong non-IID conditions ($\alpha=0.05$), achieving approximately 2 percentage points higher server-side accuracy while reducing communication costs by over 50%. On the client side, SCARLET’s accuracy closely followed the peak performance of Selective-FD (remaining within 1 percentage point of it), significantly outperforming all other baselines. This highlights the effectiveness of SCARLET’s caching mechanism and adaptive aggregation in mitigating challenges posed by heterogeneous data distributions.
TABLE V: Per-Round Uplink and Downlink Communication Costs.
Method
Uplink Cost
[MB/round]
Downlink Cost
[MB/round]
SCARLET (ours) 1.37 $\pm$ 0.28 (max: 4.30) 2.97 $\pm$ 0.29 (max: 6.32)
DS-FL [dsfl] 4.80 $\pm$ 0.00 (max: 4.80) 5.60 $\pm$ 0.00 (max: 5.60)
COMET [comet] 4.80 $\pm$ 0.00 (max: 4.80) 5.60 $\pm$ 0.00 (max: 5.60)
CFD [cfd] 1.60 $\pm$ 0.00 (max: 1.60) 5.60 $\pm$ 0.00 (max: 5.60)
Selective-FD [selectivefd] 3.88 $\pm$ 0.04 (max: 4.02) 5.52 $\pm$ 0.02 (max: 5.58)
Figure 9: Per-round uplink and downlink communication costs for SCARLET. This figure visualizes the dynamic costs summarized as averages in Table V. Unlike methods with static costs, SCARLET’s costs fluctuate dynamically, especially in the initial rounds, due to the soft-label caching mechanism. The costs stabilize as the random selection of public data staggers the timing of cache updates across individual samples, consistently maintaining a lower burden on the uplink.
To offer a more granular analysis of this communication efficiency for the CIFAR-10 dataset under a strong non-IID setting ($\alpha=0.05$), Table V details the separated uplink and downlink costs, while Fig. 9 visualizes the per-round dynamics that produce these averages for SCARLET. This breakdown is critical for practical deployments where uplink bandwidth is often a bottleneck. As shown, SCARLET’s advantage is particularly pronounced in the uplink, where the soft-label caching mechanism reduces the average cost by approximately 71% compared to DS-FL and COMET, as clients only transmit uncached data. This targeted reduction of the more critical uplink burden, while still maintaining the lowest downlink cost, underscores SCARLET’s robustness for real-world federated networks with asymmetric bandwidth.
@@
Under moderate non-IID conditions ($\alpha=0.3$), SCARLET maintained server-side accuracy comparable to the best-performing method, Selective-FD, while achieving over 50% lower communication cost. However, Selective-FD demonstrated superior client-side convergence speed and final accuracy, underscoring its strength in personalized learning scenarios. In contrast, DS-FL struggled to achieve high accuracy under these conditions, regardless of how the temperature $T$ in its ERA mechanism was tuned. This limitation can be traced to the presence of aggregated soft-labels with inherently low entropy. When $T$ is reduced excessively, these low-entropy soft-labels approach a one-hot representation, diminishing the effectiveness of the distillation process. Conversely, increasing $T$ too much causes a sudden spike in entropy for certain soft-labels, exacerbating inconsistencies in entropy reduction across public datasets. These issues result in both lower accuracy and performance degradation over time, likely due to overfitting.
@@
(a) CIFAR-100 (Server)
(b) CIFAR-100 (Client)
(c) Tiny ImageNet (Server)
(d) Tiny ImageNet (Client)
Figure 10: Test accuracy vs. communication cost on more complex datasets, CIFAR-100 and Tiny ImageNet. The results confirm the scalability and robustness of SCARLET. Even in these more challenging settings with a larger number of classes, SCARLET consistently maintains the highest communication efficiency, achieving competitive accuracy while requiring less communication than other methods. This demonstrates that the benefits of soft-label caching and Enhanced ERA generalize well beyond simple datasets.
Fig. 10 presents the results for CIFAR-100 and Tiny ImageNet, demonstrating the scalability and robustness of SCARLET. The soft-label caching mechanism and Enhanced ERA consistently achieved high communication efficiency and competitive accuracy across datasets with more complex label distributions and greater heterogeneity. On the server side, SCARLET achieves test accuracy comparable to the top-performing baselines (Selective-FD and COMET) while reducing cumulative communication costs by approximately 70%. On the client side, while Selective-FD retains a slight edge in final accuracy, SCARLET consistently ranks as a close second, delivering competitive performance that significantly outperforms the remaining baselines.
In contrast, DS-FL exhibits significantly degraded performance on these datasets. This failure highlights the inherent limitations of the conventional ERA mechanism: specifically, its sensitivity to the larger number of classes ($N$) and the practical infeasibility of finding a single optimal temperature $T$ capable of effectively regularizing the aggregated soft-labels.
@@
IV-B3 Effectiveness of the Soft-Label Caching Mechanism in Other SOTA Methods
@@
The soft-label caching mechanism, a core component of SCARLET, is designed to alleviate communication inefficiencies in distillation-based FL. Since such inefficiencies are not specific to SCARLET, the mechanism can also be applied to other SOTA methods to improve their communication efficiency. To evaluate its broader impact, we integrated the caching mechanism into three representative FL methods: CFD, COMET, and Selective-FD. DS-FL was excluded from this evaluation, as it becomes conceptually similar to SCARLET when caching is introduced. For all methods, we used a conservative cache duration of $D=25$ to balance communication reduction with the risk of overfitting to stale predictions. This evaluation was performed on CIFAR-10 under strong non-IID conditions ($\alpha=0.05$). The results presented in this section were averaged across three random seeds.
(a) CFD (Server)
(b) CFD (Client)
(c) COMET (Server)
(d) COMET (Client)
(e) Selective-FD (Server)
(f) Selective-FD (Client)
Figure 11: Impact of integrating SCARLET’s soft-label caching mechanism into other state-of-the-art methods on CIFAR-10 ($\alpha=0.05$). The results demonstrate that our caching mechanism is a modular and widely applicable component. When applied to CFD, COMET, and Selective-FD, it reduces their communication costs by approximately 50% with minimal or no degradation in server-side test accuracy. This highlights its potential as a general-purpose enhancement for distillation-based FL frameworks.
@@
Fig. 11 presents the server-side and client-side test accuracy versus cumulative communication cost for each method with the soft-label caching mechanism applied. The results show that incorporating the caching mechanism reduced server-side communication costs by approximately 50% across all three methods, without substantial loss in target model accuracy. Notably, CFD and Selective-FD maintained server-side accuracy levels comparable to their original implementations, indicating that the mechanism imposes minimal performance overhead. On the client side, although communication reductions were less dramatic due to the quicker convergence of personalized models, notable gains were still observed. Selective-FD, in particular, preserved strong personalized accuracy while benefiting from improved communication efficiency. COMET showed a notable client-side accuracy degradation in later training rounds. This is likely attributed to a fundamental conflict between COMET’s dynamic mechanism of selecting the optimal clustered teacher every round and the soft-label caching mechanism’s static nature of reusing a past teacher for $D$ rounds.
@@
These findings demonstrate the generalizability and effectiveness of the soft-label caching mechanism. Its seamless integration into a variety of distillation-based FL methods—including CFD, COMET, and Selective-FD—highlights its potential as a practical, drop-in component for enhancing communication efficiency. By incorporating this mechanism, existing FL frameworks can achieve significant server-side communication reductions. They can also maintain strong global and personalized model performance when the caching mechanism does not conflict with the method’s core design, as seen in CFD and Selective-FD. This contributes to the scalability and practicality of FL systems.
IV-B4 Ablation Study of Cache Duration